Serge’s outgoing 115-kilometer bike involved riding uphill for 2 hours and on level ground for 3
hours. On the return trip, on a different 135-kilometer trail, he rode downhill for 2 hours and on
level ground for 3 hours. Determine Serge’s three speeds if his speed on level ground is the
average of his uphill and downhill speeds.

Answer by greenestamps(13200)



Given that his speed on level ground is the average of his uphill and downhill speeds, let the three speeds be

x+y downhill
x on level ground
x-y uphill

2 hours uphill and 3 hours on level ground covered 115 km:
2(x-y)+3(x) = 115

2 hours downhill and 3 hours on level ground covered 135 km:
2(x+y)+3(x) = 135

Solve the pair of equations using basic algebra.

2x-2y+3x = 115
5x-2y = 115

2x+2y+3x = 135
5x+2y = 135

Comparing the two equations (i.e., subtracting one from the other)....

4x = 20
x = 5

Plug x=5 into one of the equations to solve for y; then use your x and y values to find the three speeds.